OpenStudy (anonymous):

Help!!!! I'll medal and fan! Solve the equation: 1/5 + w = 7 1/3

OpenStudy (anonymous):

1/5 + w = 7 1/3 w = - 3/15 + 7 5/15 w = 7 2/15.

OpenStudy (anonymous):

Thanks